# Modeling the formation of N2 and CH4 frost on Pluto’s slopes

### v1

@@ 3.3 Comparison to New Horizons observations
Figure 8: Comparisons between frost observations on Ralph/MVIC color images in the Cthulhu region reported by Bertrand et al. (2020) at Venera Terra (A), Virgil Fossae (C), Beatrice Fossae (E), and the Southern part of the region (G) and the model’s outputs (right) assuming an albedo of 0.69 for methane ice to achieve better agreement with the observations. For the performance of the baseline model (albedo of 0.67), see Figure 9. Perennial accumulation occurs on South-facing slopes (not shown).
@@
The seasonal evolution of methane frost predicted by our model is consistent with observations reported by Earle et al. (2022) and Bertrand et al. (2020) (Figures 6, 8). At high latitudes (above 40°), CH4 is predicted to be present on both north- and South-facing slopes, consistent with the detections of Earle et al. (2022) (Figures 6a,b; Figure 8b). Thick N2 frost is also expected (Figure 6a), in agreement with the observed mixtures of N2 and CH4 on these slopes.
@@
At lower latitudes (below 20°), CH4 frost is predicted to be always present on South-facing slopes, possibly diluted with N2 ice that begins to condense (Figures 6b to h; Figures 8b,d,f,g). However, neither Earle et al. (2022) nor Bertrand et al. (2020) detected frost on these South-facing slopes, although the very low illumination of these surfaces (below 0.3 W m-2, see Figure S8c of Bertrand et al. (2020)) may have prevented detection. At 15°N, methane frost is predicted to persist on the steepest North-facing slopes (at least 30°according to our model, Figures 6d, f; Figures 8d,f)) and to be sublimating. This is consistent with the reddish patterns observed next to ice deposits on such slopes, which were interpreted by Bertrand et al. (2020) as bedrock exposures indicating ongoing frost sublimation.
The exact timing of frost sublimation remains highly sensitive to surface properties and local conditions (e.g., lower thermal inertia, higher slope angle), which can promote frost stability and explain the observed variability in frost persistence (Figure 7). Notably, it appears that using a CH4 ice albedo of 0.69 leads to a better agreement than using the baseline value of 0.67 to explain the persistence of these frosts (Figures 8, 9 ). In section 3.1 we reported that frost was typically found on slopes of 12.8° $\pm$ 8° at 1-$\sigma$, but this value may be underestimated due to the limited resolution of the topographic dataset, which likely smooths slope gradients and reduces the apparent slope angles where frost is found.
@@
At the equator, CH4 frost has been observed on parts of North-facing slopes at Beatrice Fossae (Figure 8e), while our model predicts that these deposits should have disappeared by the 2000s (Figure 8f). The observed frosts are adjacent to reddish, frost-free terrains, again suggesting that local conditions strongly control frost persistence. Schenk et al. (2018) (see their Figure 23) showed that the fracture wall scarps have typical slope angle between 35° and 45°, and can locally reach up to 55°. Assuming a slope angle of 40°, our model simulates perennial accumulation of CH4 frost at this latitude (not shown).
@@ 4 Discussions
4.1 Effect of slope microclimates on Pluto’s volatiles cycles
@@
Using a similar sub-grid slope parameterization in the Mars Planetary Climate Model, Lange et al. (2023) showed that slope microclimates exert little influence on the global seasonal cycles of CO2 and H2O. This is primarily because sloped terrains (here defined as surfaces with a slope angle greater than 5°) account for only about 5% of Mars’ total area, so frost and ice on flat regions dominate the volatile cycles. In contrast, nearly 30% of Pluto’s observed surface is non-flat (Figure 3), raising the question of how N2 and CH4 condensation on slopes might affect Pluto’s global climate. To quantify this effect, we run two simulations: one where the atmosphere interacts only with flat sub-grid surfaces (all slope cover-fractions $\delta_{i}$ are set to 0 in Eqs. 9 and 10, called the ’smooth simulation’), and one where $\delta_{i}$ values in every grid cell are set to those presented in Figure 3 (the ’rough simulation’). The comparison between these simulations shows that slope microclimates do not significantly affect the N2 cycle. The rough simulation yields a total pressure of 1.14 Pa in 2015, versus 1.2 Pa for the smooth simulation. As shown in Figure 6, N2 frost begins to form on South-facing slopes rather than on flat surfaces, affecting the amount of N2 present in the atmosphere. Thick N2 frost, notably thicker than on flat surfaces, can form by the time of New Horizons flyby (Figure 6c), sometimes perennially, thereby reducing atmospheric N2 and total pressure. During Southern summer (Ls = 270°), the total pressure is higher in the rough simulation than in the smooth simulation. As highlighted in Figure 6, N2 frost on South-facing slopes sublimes, replenishing the atmosphere and explaining the increased surface pressure compared to the smooth case where N2 frost begins forming at the surface. This comparison indicates that, although Pluto’s surface is more strongly sloped than Mars and N2–CH4 frost can develop on inclined terrains where it cannot on flats, slope microclimates still appear to exert only a limited influence on Pluto’s N2 seasonal cycle. Note that in our simulations, we adopted a uniform slope distribution within each model grid cell to avoid the computational constraints discussed in Section 2.3.1. Actually, certain regions—especially those covered by thick, perennial N2 ice—are likely smoother than this idealization. In addition, the Pluto VTM treats atmospheric circulation and volatile transport in a simplified manner. Hence, a robust quantitative evaluation of slope‐driven microclimates on Pluto’s seasonal cycles will require a complete topographic map combined with simulations using the Pluto Global Climate Model, which we defer to future work.
@@
The methane cycle diagnostic is more complex. The perennial accumulation of CH4 frost on South-facing slopes and the presence of thicker frost on North-facing slopes compared to flat surfaces suggest that the rough simulation should exhibit lower atmospheric CH4 volume mixing ratios than the smooth case. However, the different total pressures between the simulations bias the results. Due to lower pressure in the rough simulation during Northern summer, methane is less diluted, resulting in higher atmospheric CH4 volume mixing ratios. Conversely, higher pressure during the Southern summer leads to greater methane dilution and lower mixing ratios compared to the smooth simulation. Since pressure variations also impact the methane cycle and the amount of methane condensing and subliming, this bias is not straightforward to correct and prevents us from drawing definitive conclusions about slope microclimate effects on the methane cycle.
@@ 4.2 Evolution of the Cthulhu region in the coming decades
As illustrated in Figures 6, 8, CH4 and N2 frosts are expected to disappear completely from slopes and flat terrains between 40° and 0° N in the Cthulhu region by around 2030. N2 frosts should instead begin to accumulate on the permanent CH4 ice deposits located on South-facing slopes. Conversely, little to no N2 or CH4 frost (thickness << 1 mm, Figure 6 ) is expected to form on the flat terrains of Cthulhu. As a result, this region should retain its dark albedo as mapped by Hofgartner et al. (2023) over the coming decades to centuries.
In the Southern hemisphere, N2 frosts a few tens of centimeters thick and a thin layer of methane frost (<< 0.1 mm) are predicted to form by 2100. Although our estimates of frost amounts remain sensitive to the model parameters and simplifications (see sections 2.3.1, 2.3.3, and 3.2.2), we can reasonably anticipate that the Northern mid-latitudes will continue to sublimate and appear darker, whereas the Southern hemisphere will gradually brighten.
@@ 5 Conclusions
In this paper, we analyze the presence of frost on slopes in the Cthulhu region of Pluto, previously thought to be frost-free due to its dark surface, and develop a parameterization in the Pluto Volatile Transport Model to explain their presence and study their stability. The main conclusions of our investigation are:
  • CH4 frost is widespread on North-facing slopes of the Cthulhu region (Figure 4), generally on steep slopes (mean slope angle 12.8° $\pm$ 8° at 1-$\sigma$, Figure 5). No frost is observed on South-facing slopes, but this may be a consequence of their very low insolation (¡0.3 W m-2) during the period of Pluto’s orbit considered, which could prevent detection.
@@
  • The Pluto Volatile Transport Model, coupled with the sub-grid slope parameterization (Figure 2), shows that frost on North-facing slopes corresponds to seasonal frost currently sublimating at the time of the New Horizons observations, in good agreement with the data (Figures 6, 8). Sublimation rates and frost persistence throughout Pluto’s orbit are highly sensitive to surface properties such as ice albedo, surface and bedrock thermal inertia, and slope angle (Figure 7).
@@
  • N2 frost deposits can form on both north- and South-facing slopes during winter in the Cthulhu region, whereas they are not expected to develop on flat surfaces (Figure 6). By the time of the New Horizons flyby, N2 frost on North-facing slopes should already have sublimated, while accumulation is expected to begin on South-facing slopes, reaching thicknesses of a few centimeters (Figure 6).
@@
  • Perennial CH4 frost deposits are expected to form on South-facing slopes (Figures 6, 8), as sublimation during Southern summer is mitigated by the higher CH4 volume mixing ratio prevailing during this season.
@@
  • Despite Pluto’s abundant slopes and the ability of frost to form—either seasonally or perennially—on inclined terrains where it cannot accumulate on flat surfaces, slope microclimates appear to play only a minor role in Pluto’s global volatile cycles (section 4.1). Conversely, the global volatile cycle strongly shapes local slope microclimates by controlling the amount of condensable material available for frost formation. Confirming the interactions between local slope-microclimates and the large-scale volatile cycles will require both a complete climate model and a detailed map of Pluto’s global topography.
@@
  • By around 2030, CH4 and N2 frosts are expected to disappear from slopes and flat terrains between 0° and 40°N in the Cthulhu region, while N2 frost will start accumulating on South-facing CH4 ice perenial deposits. Hence, the Cthulhu region should conserve its dark appearance during the next centuries. In the Southern hemisphere, N2 frosts up to a few tens of centimeters and a thin CH4 layer are predicted by 2100, suggesting that northern mid-latitudes will remain dark while the south gradually brightens (section 4.2).
@@
While the sub-grid slope model developed in this paper has so far been applied only to Mars and Pluto, the generality of our parameterization makes it suitable for studying volatile condensation and sublimation on other airless bodies or planets with thin atmospheres. In particular, coupling this parameterization with the Triton Volatile Transport Model (Bertrand et al., 2022) would allow investigation of how ice deposits form on slopes, and potentially how slopes affect seasonal volatile cycles, thus generalizing the conclusions drawn here.
@@ Acknowledgements.
LL’s research was supported by an appointment to the NASA Postdoctoral Program at the Jet Propulsion Laboratory, administered by Oak Ridge Associated Universities under contract with NASA. © 2025. All rights reserved. L.A. Young’s was funed by NASA grant NFDAP 80NSSC23K0666. Other authors acknowledge the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-23-CE49-0006 (project SHERPAS). The model used in this study can be retrieve at: https://trac.lmd.jussieu.fr/Planeto/browser/trunk/LMDZ.PLUTO. Pluto’s topographic data used is this study are from Schenk et al. (2018). CH4 band depth map and phase index retrieved from the LEISA instrument are from Schmitt et al. (2017) and Gabasova et al. (2021). Albedo maps are taken from (Hofgartner et al., 2023). Data files used to generate the figures in this analysis are available in a public repository, see Lange et al. (2025).
Appendix A Compairson between New-Horizons observations and model’s outputs using a CH4 ice albedo of 0.67.
Figure 9: Comparisons between frost observations on Ralph/MVIC color images in the Cthulhu region reported by Bertrand et al. (2020) at Venera Terra (A), Virgil Fossae (C), Beatrice Fossae (E), and the Southern part of the region (G) and the model’s outputs (right) assuming an albedo of 0.67 for methane ice to achieve better agreement with the observations.

### v2

@@ 3.3 Comparison to New Horizons observations
Figure 8: Comparisons between frost observations on Ralph/MVIC color images in the Cthulhu region reported by Bertrand et al. (2020) at Venera Terra (A), Virgil Fossae (C), Beatrice Fossae (E), and the southern part of the region (G) and the model outputs (right) assuming an albedo of 0.69 for methane ice to achieve better agreement with the observations. For the performance of the baseline model (albedo of 0.67); see Fig. 9. Perennial accumulation occurs on south-facing slopes (not shown). Perennial accumulation occurs on south-facing slopes (not shown). Our simulations predict that the CH4 frost deposits (white patches in panels A, C, E, G) observed on north-facing slopes correspond to seasonal CH4 frost that gradually sublimates and should disappear shortly after the New Horizons flyby. No curves are shown for south-facing slopes, as no CH4 frost is visible in panels A, C, E, G due to shadowing. These slopes also display only perennial accumulation, without seasonal variability (see Fig. 6).
@@
The seasonal evolution of methane frost predicted by our model is consistent with observations reported by Earle et al. (2022) and Bertrand et al. (2020, see our Figs. 6 and 8). At high latitudes (above 40°), CH4 is predicted to be present on both north and south-facing slopes, consistent with the detections of Earle et al. (2022, see our Figs. 6a, 6b, and 8b). Thick N2 frost is also expected (Fig. 6a), in agreement with the observed mixtures of N2 and CH4 on these slopes.
@@
At lower latitudes (below 20°), CH4 frost is predicted to be always present on south-facing slopes, possibly diluted with N2 ice that begins to condense (Figs. 6b to 6h and Figs. 8b, 8d, 8f, and 8g). Neither Earle et al. (2022) nor Bertrand et al. (2020) detected frost on these south-facing slopes, however, although the very low illumination of these surfaces (below 0.3 W m-2; see Fig. S8c of Bertrand et al. 2020) might have prevented detection. At 15°N, methane frost is predicted to persist on the steepest north-facing slopes (at least 30° according to our model; Figs. 6d, 6f, 8d, and 8f) and to be sublimating. This is consistent with the reddish patterns observed next to ice deposits on such slopes, which were interpreted by Bertrand et al. (2020) as bedrock exposures indicating ongoing frost sublimation.
[…]
@@
At the equator, CH4 frost has been observed on parts of north-facing slopes at Beatrice Fossae (Fig. 8e), while our model predicts that these deposits should have disappeared by the 2000s (Fig. 8f). The observed frosts are adjacent to reddish, frost-free terrains, again suggesting that local conditions strongly control frost persistence. Schenk et al. (2018, see their Fig. 23) showed that the fracture wall scarps have typical slope angle between 35° and 45°, and can locally reach up to 55°. Assuming a slope angle of 40°, our model simulates perennial accumulation of CH4 frost at this latitude (not shown).
@@ 4 Discussions
4.1 Effect of the slope microclimates on the volatile cycles of Pluto
@@
Using a similar subgrid-slope parameterization in the Mars Planetary Climate Model, Lange et al. (2023) showed that slope microclimates exert little influence on the global seasonal cycles of CO2 and H2O. This is primarily because sloped terrains (here defined as surfaces with a slope angle greater than 5°) account for only about 5% of Mars’ total area, so frost and ice on flat regions dominate the volatile cycles. In contrast, nearly 30% of the observed Pluto surface is non-flat (Fig. 3), raising the question of how N2 and CH4 condensation on slopes might affect the global climate on Pluto. To quantify this effect, we run two simulations: one where the atmosphere interacts only with flat subgrid surfaces (all slope cover-fractions $\delta_{i}$ are set to 0 in Eqs. 9 and 10, called the smooth simulation), and one where $\delta_{i}$ values in every grid cell are set to those presented in Fig. 3 (the ’rough simulation’). The comparison between these simulations shows that slope microclimates do not significantly affect the N2 cycle. The rough simulation yields a total pressure of 1.14 Pa in 2015, versus 1.2 Pa for the smooth simulation. As shown in Fig. 6, N2 frost begins to form on south-facing slopes rather than on flat surfaces, affecting the amount of N2 present in the atmosphere. Thick N2 frost, notably thicker than on flat surfaces, can form by the time of New Horizons flyby (Fig. 6c), sometimes perennially, thereby reducing atmospheric N2 and total pressure. During southern summer (Ls = 270°), the total pressure is higher in the rough simulation than in the smooth simulation. As highlighted in Fig. 6, N2 frost on south-facing slopes sublimes, replenishing the atmosphere and explaining the increased surface pressure compared to the smooth case where N2 frost begins forming at the surface. This comparison indicates that, although the Pluto surface is more strongly sloped than Mars and N2–CH4 frost can develop on inclined terrains where it cannot on flats, slope microclimates still appear to exert only a limited influence on the N2 seasonal cycle on Pluto. Note that in our simulations, we adopted a uniform slope distribution within each model grid cell to avoid the computational constraints discussed in Sect. 2.3.1. Actually, certain regions, especially those covered by thick, perennial N2 ice, are likely smoother than this idealization. In addition, the Pluto VTM treats atmospheric circulation and volatile transport in a simplified manner. Hence, a robust quantitative evaluation of slope‐driven microclimates on the seasonal cycles on Pluto will require a complete topographic map combined with simulations using the Pluto Global Climate Model, which we defer to future work.
@@
The methane cycle diagnostic is more complex. The perennial accumulation of CH4 frost on south-facing slopes and thicker frost on north-facing slopes compared to flat surfaces suggest that the rough simulation should exhibit lower atmospheric CH4 volume-mixing ratios than the smooth case. The different total pressures between the simulations bias the results, however. Due to lower pressure in the rough simulation during northern summer, methane is less diluted, resulting in higher atmospheric CH4 volume-mixing ratios. Conversely, higher pressure during the southern summer leads to greater methane dilution and lower mixing ratios compared to the smooth simulation. Since pressure variations also impact the methane cycle and the amount of methane condensing and subliming, this bias is not straightforward to correct and prevents us from drawing definitive conclusions about slope microclimate effects on the methane cycle.
@@ 4.2 Evolution of the Cthulhu region in the coming decades
As illustrated in Figs. 6 and 8, CH4 and N2 frosts are expected to disappear completely from slopes and flat terrains between 40° and 0° N in the Cthulhu region by around 2030. N2 frosts should instead begin to accumulate on the permanent CH4 ice deposits located on south-facing slopes. Conversely, little to no N2 or CH4 frost (thickness << 1 mm; Fig. 6 ) is expected to form on the flat terrains of Cthulhu. As a result, this region should retain its dark albedo as mapped by Hofgartner et al. (2023) over the coming decades to centuries.
In the southern hemisphere, N2 frosts a few tens of centimeters thick and a thin layer of methane frost (<< 0.1 mm) are predicted to form by 2100. Although our estimates of frost amounts remain sensitive to the model parameters and simplifications (see Sects. 2.3.1, 2.3.3, and 3.2.2), we can reasonably anticipate that the northern mid-latitudes will continue to sublimate and appear darker, whereas the southern hemisphere will gradually brighten.
@@ 5 Conclusions
We analyzed the frost on the slopes in the Cthulhu region of Pluto, which was previously thought to be frost-free because its surface is dark, and we developed a parameterization in the Pluto Volatile Transport Model to explain this and study the stability of the frost. The main conclusions of our investigation are listed below.
[…]
@@
  • The volatile transport model for Pluto, coupled with the subgrid-slope parameterization (Fig. 2), showed that frost on north-facing slopes corresponds to seasonal frost that currently sublimates at the time of the New Horizons observations. This agrees well with the data (Figs. 6 and 8). The sublimation rates and frost persistence throughout the Pluto orbit are highly sensitive to surface properties such as the ice albedo, the surface, the bedrock thermal inertia, and the slope angle (Fig. 7).
@@
  • N2 frost deposits can form on north- and south-facing slopes during winter in the Cthulhu region, whereas they are not expected to develop on flat surfaces (Fig. 6). At the time of the New Horizons flyby, N2 frost on north-facing slopes should already have sublimated, while accumulation is expected to begin on south-facing slopes and reach thicknesses of a few centimeters (Fig. 6).
@@
  • Perennial CH4 frost deposits are expected to form on south-facing slopes (Figs. 6 and 8), as sublimation during southern summer is mitigated by the higher CH4 volume-mixing ratio that prevails during this season.
@@
  • Despite the abundant slopes on Pluto and the ability of frost to form (either seasonally or perennially) on inclined terrains where it cannot accumulate on flat surfaces, the slope microclimates appear to play only a minor role in the global volatile cycles on Pluto (Sect. 4.1). Conversely, the global volatile cycle strongly shapes local slope microclimates by controlling the amount of condensable material available for the formation of frost. A confirmation of the interactions between local slope-microclimates and the large-scale volatile cycles will require a complete climate model and a detailed map of the global topography of Pluto.
@@
  • By about 2030, the CH4 and N2 frosts are expected to disappear from slopes and flat terrains between 0° and 40°N in the Cthulhu region, while N2 frost will start to accumulate on south-facing perennial deposits of CH4 ice. Hence, the Cthulhu region is expected to conserve its dark appearance during the next centuries. In the southern hemisphere, N2 frosts up to a few dozen centimeters and a thin CH4 layer are predicted by 2100, suggesting that the northern mid-latitudes will remain dark, while the south brightens gradually (Sect. 4.2).
  • Our model generally predicts the seasonal formation and disappearance of CH4 frost within the Cthulhu region. In particular, it reproduces the frost on north-facing slopes down to about 10°N at the time of the New Horizons flyby. These discrepancies likely stem from the model constraint, which limits slope angles to 30°, whereas local slopes can reach up to 40° along fault scarps and crater walls. When the maximum slope angle was increased in sensitivity tests, the modeled frost distribution became consistent with the observations. Future high-resolution observations might confirm perennial frost on south-facing slopes. On the other hand, new James Webb Space Telescope observations of the northern and southern Pluto hemispheres could test our prediction that the northern frosts are progressively vanishing as the southern hemisphere enters winter, and this might help us to constrain key model parameters, such as the optical properties of the ice and the thermal inertia of the substrate (e.g., Bertrand et al., 2025), which are currently poorly constrained.
@@
While the subgrid-slope model we developed was only applied to Mars and Pluto so far, the generality of our parameterization makes it suitable for studying volatile condensation and sublimation on other airless bodies or planets with thin atmospheres. In particular, coupling this parameterization with the volatile transport model for Triton (Bertrand et al., 2022) would allow us to investigate the formation of ice deposits on slopes, and potentially, the effect of slopes on seasonal volatile cycles. This would allow us to generalize the conclusions drawn here.
@@ Acknowledgements.
LL’s research was supported by an appointment to the NASA Postdoctoral Program administered by Oak Ridge Associated Universities under contract with NASA at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). © 2025. All rights reserved. L.A. Young’s was funded by NASA grant NFDAP 80NSSC23K0666. Other authors acknowledge the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-23-CE49-0006 (project SHERPAS). The model used in this study can be retrieved at: https://trac.lmd.jussieu.fr/Planeto/browser/trunk/LMDZ.PLUTO. The topographic Pluto data we used are taken from Schenk et al. (2018). CH4 band depth map and phase index retrieved from the LEISA instrument are from Schmitt et al. (2017) and Gabasova et al. (2021). Albedo maps are taken from (Hofgartner et al., 2023). Data files used to generate the figures in this analysis are available in a public repository, see Lange et al. (2025).
Appendix A Comparison between New-Horizons observations and model outputs using a CH4 ice albedo of 0.67
Figure 9: Comparisons between frost observations on Ralph/MVIC color images in the Cthulhu region reported by Bertrand et al. (2020) at Venera Terra (A), Virgil Fossae (C), Beatrice Fossae (E), and the southern part of the region (G) and the model outputs (right) assuming a baseline albedo of 0.67 for methane ice. This figure is the analog of Fig. 8, with all model parameters kept identical except that the albedo of CH4 ice was set to its baseline value of 0.67 instead of 0.69.